OpenStudy (anonymous):

Solve using substitution: y=x-2 x+y=10

OpenStudy (anonymous):

x+y=10 x+x+2=10 2x+2=10 2x=8 x=4

OpenStudy (anonymous):

y=4-2 y=2

OpenStudy (anonymous):

oops minus 2 x+x-2=10 +2=12 2x-2=10 2x=12 x=6

OpenStudy (anonymous):

y=6-2 y=4

OpenStudy (anonymous):

theres the right answer sorry

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

x=6 y=4

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!! :D :D

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

thanks for medal ;)

OpenStudy (anonymous):

Your Welcome!! :)